OpenStudy (anonymous):

Medal will be given! Lisa has only nickels and dimes in her money box. She knows that she has less than $15 in the box. Let x represent the number of nickels in the box and y represent the number of dimes in the box. @texaschic101 @SolomonZelman @ganeshie8

OpenStudy (anonymous):

Draw a dashed line to represent the graph of 5x + 10y = 1500, and shade the portion below the line for positive values of x and y. Draw a dashed line to represent the graph of 5x + 10y = 1500, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 1500, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x – 5y = 1500, and shade the portion below the line for positive values of x and y.

OpenStudy (anonymous):

I chose A

OpenStudy (solomonzelman):

d - dimes n - nickels \(\normalsize\color{blue}{ 10d + 5n < 1500}\) We need to know the number of coins in the box though.

OpenStudy (solomonzelman):

So that we can have an equation to represent the overall amount in dollars, and the number of coins.

OpenStudy (texaschic101):

if it's less then, wouldn't it be shaded below the line

OpenStudy (anonymous):

yes it would @texaschic101

OpenStudy (solomonzelman):

should be

OpenStudy (solomonzelman):

And below a DASHED line

OpenStudy (texaschic101):

then I agree...I think it is A

OpenStudy (solomonzelman):

Yes, A.. Agreed

OpenStudy (solomonzelman):

I misread the question at first, sorry -:(

OpenStudy (texaschic101):

it;s unanimous :) It is A

OpenStudy (anonymous):

Can you check 5 more?

OpenStudy (anonymous):

The graph of an equation is shown below: line joining ordered pairs negative 3,2 and 1, negative 2 Based on the graph, which of the following represents a solution to the equation? (–2,–3) (3, 1) (1, 3) (3, 2)